Macroeconomic

Mathematical Note 8. An island economy produces only fish and crabs. Quantities Fish Crabs Prices Fish Crabs 2016 2017 1,100 tons 525 tons 1,000 tons 500 tons 2016 $20 a ton $10 a ton Calculate the islands chained-dollar real GDP in 2017 expressed in 2016 dollars. 2017 $30 a ton $8 a ton
Answer #1

2016 quantities at 2016 prices = 1000 x 20 + 500 x 10 = $ 25000

2017 quantities at 2016 prices = 1100 x 20 + 525 x 10 = $ 27250

2016 quantities at 2017 prices = 1000 x 30 + 500 x 8 = $ 34000

2017 quantities at 2017 prices = 1100 x 30 + 525 x 8 = $ 37200

growth rate of GDP at 2016 prices = (27250-25000)/25000 x 100 = 9%

growth rate of GDP at 2017 prices = (37200-34000)/34000 x 100 = 9.41%

Average growth rate = 9.206%

Real GDP in 2017 = 25000 x (1 + 9.206%) = $ 27301.47